Water is flowing through a horizontal pipe with an
Answer #1

using equation of continuity

V1A1= V2A2

V1 \pi( 4)^2 = 27 \pi(1.875)^2

V1= 5.93 m/s apprx

b) use bernoulli's

P1 + 0.5 \rhov1^2 = p2 + 0.5 \rho​v2^2

670 x 10^3 + 0.5 ( 1x 10^3) ( 5.93)^2 = P2 +  0.5 ( 1x 10^3) ( 27 )^2

670000 + 17582.45 = P2 + 364500

P2= 323 kPa apoprx